# Improved Treatment of Bosonic Dark Matter Dynamics in Neutron Stars: Consequences and Constraints

### v1

@@ Abstract
It is conceivable that a bosonic dark matter (DM) with non-gravitational interactions with SM particles will be accumulated at the center of a neutron star (NS) and can lead to black hole formation. In contrast to previous works with a fixed NS temperature, we dynamically determine the formation of Bose-Einstein condensate (BEC) for a given set of DM parameters, namely the DM-neutron scattering cross-section ($\sigma_{\chi n}$), the thermal average of DM annihilation cross-section ($\expectationvalue{\sigma v}$) and the DM mass ($m_{\chi}$). For both non-annihilating and annihilating DM with $\expectationvalue{\sigma v}\leq 10^{-26}{~{}\rm cm^{3}~{}s^{-1}}$, the BEC forms for $m_{\chi}\lesssim 10$ TeV. In case of non-annihilating DM, observations of old NS allow $\sigma_{\chi n}\lesssim 10^{-52}~{}{\rm cm^{2}}$ for $10{~{}\rm MeV}\leq m_{\chi}\lesssim 10{~{}\rm GeV}$ (with BEC) and $\sigma_{\chi n}\lesssim 10^{-47}~{}{\rm cm^{2}}$ for $5{~{}\rm TeV}\lesssim m_{\chi}\lesssim 30{~{}\rm PeV}$ (without BEC). This analysis shows the mass window, $10{~{}\rm GeV}\lesssim m_{\chi}\lesssim 5{~{}\rm TeV}$ unconstrained, subject only to direct detection experiments. In the annihilating DM scenario, the exclusion limits on DM parameters become weaker and even vanish for typical WIMP annihilation cross-section. However, the late-time heating of the NS enables us to probe the region with $\sigma_{\chi n}\gtrsim 10^{-47}~{}{\rm cm^{2}}$, using the James Webb Space Telescope in the foreseeable future. When our results are viewed in the context of indirect searches of DM, it provides a lower limit on the $\expectationvalue{\sigma v}$, which is sensitive to the DM thermal state.
@@ 1 Introduction and Summary
The first hint of probing DM interactions in NSs comes from the “missing pulsar problem” Pfahl:2003tf ; Macquart:2010vf ; Dexter:2013xga , which questions the null observation old millisecond pulsars near the center of the Milky Way galaxy. In Ref.Bramante:2014zca , it has been pointed out that DM acquisition inside the NS can lead to the black hole (BH) formation and eventual destruction of the host star, thereby explaining the missing pulsar population. In this context, the observation of an old NS with typical age about a Gyr near the galactic center GC ; Bower:2013tva or any other DM-rich environments (like globular clusters Camilo:2005aa ; Padmanabh:2024bsz ; Brown:2018pwq ; Reynoso-Cordova:2022ojo ) can constrain the DM capture rate. It is evident that the constraint is dependent on the age of the NS as well as the DM density around it. In particular, the older the NS and the denser the local DM halo, the stronger the constraint on the DM capture rate becomes. Now, the DM capture takes place only when the gravitationally in-falling DM particles dissipate their kinetic energy via elastic scatterings with neutrons and eventually get trapped inside the NS. Consequently, the DM-neutron scattering cross-section [^1] From now on, $\sigma_{\chi n}$ represents the DM-neutron scattering cross-section. ($\sigma_{\chi n}$) get constrained from the DM capture rate for a given DM mass. In fact, the constraint on $\sigma_{\chi n}$ is expected to be stronger as well as for a wider mass range than in the terrestrial direct detection experiments due to higher density of neutrons in NSs deLavallaz:2010wp ; Raj:2017wrv ; Maity:2021fxw .
Now, with non-zero annihilation rate of DM particles inside the NS, the upper bound on $\sigma_{\chi n}$ weakens. Moreover it can become completely unconstrained from the mere existence of old NS in presence of sufficiently large DM annihilation cross-section. However, the DM annihilation to SM states confined within the NS modifies the standard evolution of NS cooling Yakovlev:2003qy , that results in the late-time heating of an old NS. Thermal emissions from the surface of an old NS, associated with this late-time heating can be detectable in current and future probes, such as the James Webb Space Telescope (JWST)JWST , Thirty Meter Telescope (TMT)TMT or Extremely Large Telescope (ELT)ELT with sensitivities in the visible and infra-red frequencies. Such a detection can again constrain $\sigma_{\chi n}$ which decides the amount of heating due to DM annihilation in the capture-annihilation equilibrium condition Kouvaris:2007ay .
@@
In the present study, we dynamically determine (in terms of parameters namely $\sigma_{\chi n}$, $\expectationvalue{\sigma v}$, $m_{\chi}$) the thermal state of bosonic DM particles inside a NS, which is not considered in most of the previous studies McDermott:2011jp ; Bramante:2013hn ; Garani:2018kkd ; Bhattacharya:2023stq ; Lu:2024kiz , as per our knowledge. In particular, we at first focus on the distinction between a BEC and a non-BEC state in the DM parameter space. Then we follow its effect on possible BH formation and the late-time heating for annihilating DM scenarios. Our main findings are summarized below.
@@
    For a BEC state, constraints on the $\sigma_{\chi n}$ derived from the observation of $10$ Gyr NS is stronger than in the non-BEC state owing to larger DM particles requiring for BH formation in the latter case. In the non-annihilating DM scenario, masses within $10~{}{\rm MeV}\lesssim m_{\chi}\lesssim 10$ GeV, are disallowed for $\sigma_{\chi n}$ values as shown in Fig.3 (red shaded region) for the BEC state. For a non BEC state, the BH formation due to self-gravitation constrains DM masses $10~{}{\rm TeV}\lesssim m_{\chi}\lesssim 30~{}{\rm PeV}$ as shown in Fig.3 (green shaded region). We note that the constraint on $\sigma_{\chi n}$ for the non-BEC state is found to be weaker significantly in our dynamical analysis than that of the Refs.McDermott:2011jp ; Garani:2018kkd . For detailed discussion on apparent differences see Sec.3.
@@
    In the non-annihilating DM scenario, the regions in the DM parameter space with $10~{}{\rm GeV}\lesssim m_{\chi}\lesssim 5~{}{\rm TeV}$ and $m_{\chi}\gtrsim 30~{}{\rm PeV}$ are not constrained from the existence of an old NS, due to BH evaporation through Hawking radiation. Therefore, the direct detection (DD) experiments remain important to probe the DM-neutron scattering cross-section for $10~{}{\rm GeV}\lesssim m_{\chi}\lesssim 5~{}{\rm TeV}$. However, for sub-GeV DM masses ($10~{}{\rm MeV}\lesssim m_{\chi}\lesssim 10~{}{\rm GeV}$), the NS observations can be complementary to the DD experiments.
@@
    In the annihilating DM scenarios, the upper bounds on the $\sigma_{\chi n}$ for BEC and non-BEC states weaken due to depletion of DM particles, signifying the survival of an old NS. We find that for canonical WIMP scenario with $\expectationvalue{\sigma v}\approx 10^{-26}{~{}\rm cm^{3}s^{-1}}$, there is no constraint in the $\sigma_{\chi n}-m_{\chi}$ plane from the existence of a $10$ Gyr old NS. However, the DM annihilation-induced late-time heating of NS has been exploited to show the parameter space for $\sigma_{\chi n}$, which can be probed by the JWST in the foreseeable future.
@@ 2.1 Dynamical setup
Formation of BEC state primarily depends on the DM capture rate, the ambient temperature of the NS and the DM annihilation rate. The deposition of DM particles into an isolated NS is facilitated by strong gravitational potential of the NS. However, the in-falling DM particles DM particles can not be trapped inside a NS without any interaction between DM and neutrons present in the NS. This stems from the fact that DM particles acquire the escape velocity of the host star, thereby escape from the interior of the star. Therefore, the capture rate depends on the scattering between DM and neutrons, given the local DM density ($\rho_{\chi}$) as follows [^2] For more accurate calculation of DM capture rate, including effects, like - relativistic energy transfer, Pauli blocking and multiple scatterings see Refs.Bell:2020jou ; Bramante:2017xlb ; Dasgupta:2019juq .Kouvaris:2007ay ; Bell:2013xk :
@@
The difference in thermalization time around $m_{\chi}\sim 1$ GeV, originates from the difference in the momentum transfer ($\delta p$) during the elastic scatterings, depending on the DM mass as $\delta p\sim(m_{\chi}m_{n})/(m_{\chi}+m_{n})$. For more detailed analysis on thermalization see Refs.Bertoni:2013bsa ; Bell:2023ysh ; Garani:2018kkd . Thermalized DM particles forms a dark core with thermal radius $r_{th}$ given by Bell:2023ysh ,
@@
where $\rho_{c}=5\times 10^{38}~{}{\rm GeV~{}cm^{-3}}$, being the typical baryon density of a NS core. This stems from the balance between gravitational force acted on DM particles due to baryon density within radius $r_{th}$ and the pressure due to thermal motions of DM particles with an approximate average velocity of $\sqrt{T_{NS}/m_{\chi}}$.
@@
It is apparent that $T_{c}$ is a dynamical quantity as it depends on the ambient number ($N_{\chi}$) of DM particles as well as the core temperature of a NS via its dependence through $r_{th}$. Therefore, we exploit the NS temperature evolution in presence of DM capture events to determine the epoch of BEC formation within the lifetime of a NS. In fact, the BEC formation is completely determined by the DM parameters, namely $\sigma_{\chi n}$, $\expectationvalue{\sigma v}$ and $m_{\chi}$.
@@
In the capture-annihilation equilibrium, the dark emissivity becomes independent of the ambient NS temperature (apparent from Eq.11), thereby $T_{NS}$ freezes out at a constant value, $T^{f}_{NS}$, signifying the late-time heating due to DM annihilation [^4] Apart from DM annihilations, there can other sources for late-time heating of a NS. For such examples see Refs.Baryakhtar:2017dbj ; Fujiwara:2023hlj ; Raj:2024kjq and references therein..
@@
Noticeably, $N_{sg}(t)$, $T_{sg}(t)$ are too functions of time, which makes our dynamical analysis essential. In terms of characteristic temperatures BEC forms when $T_{sg}<T_{NS}<T_{c}$. DM particles become self-gravitating without forming a BEC when $T_{c}<T_{NS}<T_{sg}$.
@@ 2.2 Black hole formation inside a neutron star
Here $c_{s}$ is the speed of sound inside the star. The nascent BH starts growing if $dM_{BH}/dt>0$, i.e. the accretion is larger than the evaporation rate, which in turn gives the following lower bound on the initial mass of BH with negligible Hawking radiation.
@@ 3 Results and Discussions
In this section, we simultaneously solve the evolution equations of NS temperature (Eq.7) and the yield of DM particles (Eq.10) inside the NS residing in a DM halo with $\rho_{\chi}=10^{3}~{}{\rm GeV~{}cm^{-3}}$, for given DM parameters : DM-neutron scattering cross-section ($\sigma_{\chi n}$), DM mass ($m_{\chi}$) and DM annihilation cross-section ($\expectationvalue{\sigma v}$). We set initial conditions as, $T^{i}_{NS}=1$ MeV and $N^{i}_{\chi}=0$. Consequently, we determine the thermal state of DM as well as the stability of the NS. Our goal is to find the DM parameter space allowed by observations of old NSs. As we will see, for some part of the parameter space DM particles will form either a BEC or a non-BEC state. For the subset of this parameter space, the host star transmutes to a BH. We present our results separately for annihilating and non-annihilating DM scenarios.

### v2

@@ Abstract
It is conceivable that a bosonic dark matter (DM) with non-gravitational interactions with SM particles will be accumulated at the center of a neutron star (NS) and can lead to black hole formation. In contrast to previous works with a fixed NS temperature, we dynamically determine the formation of Bose-Einstein condensate (BEC) for a given set of DM parameters, namely the DM-neutron scattering cross-section ($\sigma_{\chi n}$), the thermal average of DM annihilation cross-section ($\expectationvalue{\sigma v}$) and the DM mass ($m_{\chi}$). For both non-annihilating and annihilating DM with $\expectationvalue{\sigma v}\lesssim 10^{-26}{~{}\rm cm^{3}~{}s^{-1}}$, the BEC forms for $m_{\chi}\lesssim 10$ TeV. In case of non-annihilating DM, observations of old NS allows $\sigma_{\chi n}\lesssim 10^{-52}~{}{\rm cm^{2}}$ for $10{~{}\rm MeV}\leq m_{\chi}\lesssim 10{~{}\rm GeV}$ (with BEC) and $\sigma_{\chi n}\lesssim 10^{-47}~{}{\rm cm^{2}}$ for $5{~{}\rm TeV}\lesssim m_{\chi}\lesssim 30{~{}\rm PeV}$ (without BEC). This analysis shows that the electroweak mass window, $10{~{}\rm GeV}\lesssim m_{\chi}\lesssim 5{~{}\rm TeV}$ is essentially unconstrained by NS observations and therefore is subject only to direct detection experiments. In the annihilating DM scenario, the exclusion limits on DM parameters become weaker and even vanish for typical WIMP annihilation cross-section. However, the late-time heating of the NS enables us to probe the region with $\sigma_{\chi n}\gtrsim 10^{-47}~{}{\rm cm^{2}}$, using the James Webb Space Telescope in the foreseeable future. When our results are viewed in the context of indirect searches of DM, it provides a lower limit on the $\expectationvalue{\sigma v}$, which is sensitive to the DM thermal state.
@@ 1 Introduction and Summary
The first hint of probing DM interactions in NSs comes from the “missing pulsar problem” Pfahl:2003tf ; Macquart:2010vf ; Dexter:2013xga , which questions the null observation old millisecond pulsars near the center of the Milky Way galaxy. In Ref.Bramante:2014zca , it has been pointed out that DM acquisition inside the NS can lead to the black hole (BH) formation and eventual destruction of the host star, thereby explaining the missing pulsar population. In this context, the observation of an old NS with typical age about a Gyr near the galactic center GC ; Bower:2013tva or any other DM-rich environments (like globular clusters Camilo:2005aa ; Padmanabh:2024bsz ; Brown:2018pwq ; Reynoso-Cordova:2022ojo ) can constrain the DM capture rate. It is evident that the constraint is dependent on the age of the NS as well as the DM density around it. In particular, the older the NS and the denser the local DM halo, the stronger the constraint on the DM capture rate becomes. Now, the DM capture takes place only when the gravitationally in-falling DM particles dissipate their kinetic energy via elastic scatterings with neutrons and eventually get trapped inside the NS. Consequently, the DM-neutron scattering cross-section [^1] From now on, $\sigma_{\chi n}$ represents the DM-neutron scattering cross-section. ($\sigma_{\chi n}$) get constrained from the DM capture rate for a given DM mass. In fact, the constraint on the $\sigma_{\chi n}$ is expected to be stronger and applicable for a wider mass range than in the terrestrial direct detection experiments due to higher density of neutrons in NSs deLavallaz:2010wp ; Raj:2017wrv ; Maity:2021fxw ; Linden:2024uph .
Now, with non-zero annihilation rate of DM particles inside the NS, the upper bound on $\sigma_{\chi n}$ weakens. Moreover it can become completely unconstrained from the mere existence of old NS in presence of sufficiently large DM annihilation cross-section. However, the DM annihilation to SM states confined within the NS modifies the standard evolution of NS cooling Yakovlev:2003qy , that results in the late-time heating of an old NS. Thermal emissions from the surface of an old NS, associated with this late-time heating can be detectable in current and future probes, such as the James Webb Space Telescope (JWST)JWST , Thirty Meter Telescope (TMT)TMT or Extremely Large Telescope (ELT)ELT with sensitivities in the visible and infra-red frequencies. Such a detection can again constrain the $\sigma_{\chi n}$ which decides the amount of heating due to DM annihilation in the capture-annihilation equilibrium condition Kouvaris:2007ay .
@@
In the present study, we dynamically determine (in terms of parameters namely $\sigma_{\chi n}$, $\expectationvalue{\sigma v}$, $m_{\chi}$) the thermal state of bosonic DM particles inside a NS, which is not considered in most of the previous studies McDermott:2011jp ; Bramante:2013hn ; Garani:2018kkd ; Bhattacharya:2023stq ; Lu:2024kiz ; Liu:2024qbe ; Ray:2023auh , as per our knowledge. In particular, we at first focus on the distinction between a BEC and a non-BEC state in the DM parameter space. Then we follow its effect on possible BH formation and the late-time heating for annihilating DM scenarios. Our main findings are summarized below.
@@
    For a BEC state, constraints on the $\sigma_{\chi n}$ derived from the observation of $10$ Gyr old NS are stronger than in the non-BEC state owing to larger DM particles requiring for BH formation in the latter case. In the non-annihilating DM scenario, masses within $10~{}{\rm MeV}\lesssim m_{\chi}\lesssim 10$ GeV, are disallowed for $\sigma_{\chi n}$ values as shown in Fig.3 (red shaded region) for the BEC state. For a non BEC state, the BH formation due to self-gravitation constrains DM masses $10~{}{\rm TeV}\lesssim m_{\chi}\lesssim 30~{}{\rm PeV}$ as shown in Fig.3 (green shaded region). We note that the constraint on $\sigma_{\chi n}$ for the non-BEC state is found to be weaker significantly in our dynamical analysis than that of the Refs.McDermott:2011jp ; Garani:2018kkd . For detailed discussion on apparent differences see Sec.3.
@@
    In the non-annihilating DM scenario, the regions in the DM parameter space with $10~{}{\rm GeV}\lesssim m_{\chi}\lesssim 5~{}{\rm TeV}$ and $m_{\chi}\gtrsim 30~{}{\rm PeV}$ are not constrained from the existence of an old NS, due to BH evaporation through Hawking radiation. Therefore, the direct detection (DD) experiments remain important to probe the DM-neutron scattering cross-section for DM masses within $10~{}{\rm GeV}\lesssim m_{\chi}\lesssim 5~{}{\rm TeV}$. However, for sub-GeV DM masses ($10~{}{\rm MeV}\lesssim m_{\chi}\lesssim 10~{}{\rm GeV}$), the NS observations can be complementary to the DD experiments.
@@
    In the annihilating DM scenarios, the upper bounds on the $\sigma_{\chi n}$ for BEC and non-BEC states weaken due to depletion of DM particles, signifying the survival of an old NS. We find that for canonical WIMP scenario with $\expectationvalue{\sigma v}\approx 10^{-26}{~{}\rm cm^{3}~{}s^{-1}}$, there is no constraint in the $\sigma_{\chi n}-m_{\chi}$ plane from the existence of a $10$ Gyr old NS. However, the DM annihilation-induced late-time heating of NS has been exploited to show the parameter space for $\sigma_{\chi n}$, which can be probed by the JWST in the foreseeable future.
@@ 2.1 Dynamical setup
Formation of BEC state primarily depends on the DM capture rate, the ambient temperature of the NS and the DM annihilation rate. The deposition of DM particles into an isolated NS is facilitated by strong gravitational potential of the NS. However, the in-falling DM particles can not be trapped inside a NS without any interaction between DM and neutrons present in the NS. This stems from the fact that DM particles acquire the escape velocity of the host star, thereby escape from the interior of the star. Therefore, the capture rate depends on the scattering between DM and neutrons, given the local DM density ($\rho_{\chi}$) as follows [^2] For more accurate calculation of DM capture rate, including effects, like - relativistic energy transfer, Pauli blocking and multiple scatterings see Refs.Bell:2020jou ; Bramante:2017xlb ; Dasgupta:2019juq .Kouvaris:2007ay ; Bell:2013xk :
@@
The difference in thermalization time around $m_{\chi}\sim 1$ GeV, originates from the difference in the momentum transfer ($\delta p$) during the elastic scatterings, depending on the DM mass as $\delta p\sim(m_{\chi}m_{n})/(m_{\chi}+m_{n})$. For more detailed analysis on thermalization see Refs.Bertoni:2013bsa ; Bell:2023ysh ; Garani:2018kkd . Thermalized DM particles form a dark core with thermal radius $r_{th}$ given by Bell:2023ysh ,
@@
where $\rho_{c}=5\times 10^{38}~{}{\rm GeV~{}cm^{-3}}$, being the typical baryon density of a NS core. This stems from the balance between the gravitational force acted on DM particles due to baryon density within radius $r_{th}$ and the pressure due to thermal motions of DM particles with an approximate average velocity of $\sqrt{T_{NS}/m_{\chi}}$.
@@
$T_{c}$ determines the transition of bosons into a condensed state in which most of the particles reside in the ground state. It is apparent that $T_{c}$ is a dynamical quantity as it depends on the ambient number ($N_{\chi}$) of DM particles as well as the core temperature of a NS via its dependence through $r_{th}$. Therefore, we exploit the NS temperature evolution in presence of DM capture events to determine the epoch of BEC formation within the lifetime of a NS. In fact, the BEC formation is completely determined by the DM parameters, namely $\sigma_{\chi n}$, $\expectationvalue{\sigma v}$ and $m_{\chi}$.
@@
In the capture-annihilation equilibrium, the dark emissivity becomes independent of the ambient NS temperature (apparent from Eq.11), thereby $T_{NS}$ freezes out at a constant value, $T^{f}_{NS}$, signifying the late-time heating due to DM annihilation [^4] Apart from DM annihilations, there can other sources for late-time heating of a NS. For such examples see Refs.Baryakhtar:2017dbj ; Fujiwara:2023hlj ; Raj:2024kjq ; Ema:2024wqr and references therein..
@@
Noticeably, $N_{sg}(t)$, $T_{sg}(t)$ are too functions of time, which make our dynamical analysis essential. In terms of characteristic temperatures BEC forms when $T_{sg}<T_{NS}<T_{c}$. DM particles become self-gravitating without forming a BEC when $T_{c}<T_{NS}<T_{sg}$.
@@ 2.2 Black hole formation inside a neutron star
Here $c_{s}(=0.17c)$ is the speed of sound inside the star. The nascent BH starts growing if $dM_{BH}/dt>0$, i.e. the accretion is larger than the evaporation rate, which in turn gives the following lower bound on the initial mass of BH with negligible Hawking radiation.
@@
With this one can estimate the collapse time Singh:2022wvw , within which the small black hole acquires more baryonic matter and eventually devour the entire host star, forming a BH with a mass, equal to the NS mass. The collapse time is calculated from Eq.21 ignoring the second term as,
$t_{c}\simeq\frac{c^{3}_{s}}{4\pi\rho_{c}G^{2}M^{i}_{BH}}\simeq 40~{}{\rm years}\left(\frac{10^{42}~{}{\rm GeV}}{M^{i}_{BH}}\right).$ (22)
As is evident that $t_{c}$ depends on the initial mass of the BH, which is determined by primeval thermal state of DM particles, leading to the BH formation. Therefore, the collapse time for BEC and non-BEC cases are quite different, as apparent in the following.
$t_{c}=\begin{cases}4.6\times 10^{5}~{}{\rm years}~{}\left(\frac{m_{\chi}}{\rm GeV}\right)~{}~{}\text{for BEC BH},\\ 3.6\times 10^{2}~{}{\rm years}~{}\left(\frac{m_{\chi}}{10^{4}~{}\rm GeV}\right)^{3/2}\left(\frac{10^{-5}~{}\rm MeV}{T_{NS}}\right)^{3/2}~{}~{}\text{for non-BEC BH}.\end{cases}$ (23)
This difference becomes instrumental in distinguishing a BEC state BH from the other in the gravitational wave signal related to the binary system. We have briefly discussed the distinction between these two kinds BHs originated from collapse of NSs in Sec.4.
@@ 3 Results and Discussions
In this section, we simultaneously solve the evolution equations of NS temperature (Eq.7) and the yield of DM particles (Eq.10) inside the NS residing in a DM halo with $\rho_{\chi}=10^{3}~{}{\rm GeV~{}cm^{-3}}$, for given DM parameters : DM-neutron scattering cross-section ($\sigma_{\chi n}$), DM mass ($m_{\chi}$) and DM annihilation cross-section ($\expectationvalue{\sigma v}$). We work with initial conditions as, $T^{i}_{NS}=1$ MeV and $N^{i}_{\chi}=0$. The NS temperature becomes only a function of time within first $100$ years of its birth, therefore we start our analysis from $t=100$ years with $T_{NS}=1$ MeV, consistent with standard cooling mechanism of a NS. Subsequently, we determine the thermal state of DM as well as the stability of the NS. Our goal is to find the DM parameter space allowed by observations of old NSs. As we will see, for some part of the parameter space DM particles will form either a BEC or a non-BEC state. For the subset of this parameter space, the host star transmutes to a BH. We present our results separately for annihilating and non-annihilating DM scenarios. All the results are shown for typical NS parameters, namely $M_{NS}=1.4M_{\odot}$ and $R_{NS}=10.6$ km. Given the value of $R_{NS}$, the NS mass can vary within $(0.8-2.2)M_{\odot}$ Bramante:2023djs , which modifies the DM capture rate by less than factor of 2 as compared to our present choice of NS parameters.